3 Rosenberg sisters ages are consecutive odd integers. if the sum of their ages is 231, how old is oldest sister

Answer by stanbon(75887)

3 Rosenberg sisters ages are consecutive odd integers. if the sum of their ages is 231, how old is oldest sister
------------
1st: 2x-1
2nd: 2x+1
3rd: 2x+3
----------
Sum = 231
6x+3 = 231
6x = 228
x = 38
------
Oldest = 2x+3 = 2*38+3 = 89 years
--------------------
Cheers,
Stan H.
====================